The earth has a radius of approximately 6400 km. How many radians per second is the earth rotating at the equator? What is the speed of the earth, in kilometres per hour? If the earth rotates 2, how many kilometres has it rotated at the origin?
Answer by Alan3354(69443)

The earth has a radius of approximately 6400 km. How many radians per second is the earth rotating at the equator? What is the speed of the earth, in kilometres per hour? If the earth rotates 2, how many kilometres has it rotated at the origin?
---------------
How many radians per second is the earth rotating at the equator?
The rotation is 2pi in 24 hours.
rad/sec = 2pi/(24*60*60)
= 0.000072722... rad/sec, = 7.2722E-5, at the equator, and at every point on the earth. The radius doesn't matter, either.
----------------------
What is the speed of the earth, in kilometres per hour?
Do you mean rotational? At the equator? If so:
v = 6400*2pi/24 = ~ 1675.5 km/hour
-------
The speed of the earth in orbit is apx 18 miles/second (around the sun), or 67,000 mph.
---------------------
If the earth rotates 2, how many kilometres has it rotated at the origin?
I suppose you mean 2 rotations.
6400*2pi*2 = 25,600 pi kilometers = ~ 80425 km